How many sides does a polygon with an interior angle measure of 108 degrees have? Show your work.
Answer by Alan3354(69443)

How many sides does a polygon with an interior angle measure of 108 degrees have? Show your work.
---------
Ext angles = 180 - 108 = 72
# of sides = 360/72 = 5